### v1

# Accelerating Diffusion Transformers with Dual Feature Caching

Abstract
Diffusion Transformers (DiT) have become the dominant methods in image and video generation yet still suffer substantial computational costs. As an effective approach for DiT acceleration, feature caching methods are designed to cache the features of DiT in previous timesteps and reuse them in the next timesteps, allowing us to skip the computation in the next timesteps. However, on the one hand, aggressively reusing all the features cached in previous timesteps leads to a severe drop in generation quality. On the other hand, conservatively caching only the features in the redundant layers or tokens but still computing the important ones successfully preserves the generation quality but results in reductions in acceleration ratios. Observing such a tradeoff between generation quality and acceleration performance, this paper begins by quantitatively studying the accumulated error from cached features. Surprisingly, we find that aggressive caching does not introduce significantly more caching errors in the caching step, and the conservative feature caching can fix the error introduced by aggressive caching. Thereby, we propose a dual caching strategy that adopts aggressive and conservative caching iteratively, leading to significant acceleration and high generation quality at the same time. Besides, we further introduce a V-caching strategy for token-wise conservative caching, which is compatible with flash attention and requires no training and calibration data.
Our codes have been released in Github:
Code: https://github.com/Shenyi-Z/DuCa
1 Introduction
Diffusion Models (DMs)[16] have shown remarkable success in both image [27] and video generation [2]. Traditional DMs often use U-Net denoising structures, such as Stable Diffusion (SD)[27] and Stable Video Diffusion (SVD)[2], which have already demonstrated impressive results. Most recently, with the introduction of Diffusion Transformers (DiT)[26], the quality of visual generation has further significantly improved, propelling the development of numerous downstream applications [12]. However, the quality improvement is also accompanied by a significant increase in computational demands, reducing the inference efficiency of transformers and making their deployment in practical scenarios a pressing challenge.
Recently, feature caching has become one of the most popular techniques for diffusion acceleration. Motivated by the high similarity between features in the adjacent timesteps, aggressive feature caching caches the features computed in a freshing timestep, and then reuses them in the following several caching steps, as shown in Figure 1(b). Such a fresh-then-cache strategy enables the diffusion model to save computational costs in the caching step and thus leads to significant acceleration. However, since the similarity of features in non-adjacent timesteps decreases quickly, the error introduced by aggressive feature caching can be accumulated at an exponential speed, leading to a significant drop in generation quality, and making aggressive feature caching impractical.
Figure 1: Comparison between conservative caching (a) and aggressive caching (b): Conservative caching skips fewer layers than aggressive caching to maintain the generation quality, but also leads to lower acceleration ratios. $t$ and $t-1$ denote the refreshing and caching timesteps, respectively.
To solve this problem, recent works [8, 24, 28, 35] introduce a few computations during the caching steps in certain layers and tokens, which make the cached features in previous steps be corrected for the current caching step. These conservative feature caching methods reduce the accumulated error from feature caching but also make it difficult to achieve a higher acceleration ratio. Such a tradeoff between acceleration ratio and generation quality becomes the most essential problem for feature caching methods.
Figure 2: The caching error and acceleration ratio of three feature caching strategies. The caching error is defined as the $L2$-norm distance between $x_{t}$ computed with and without feature caching.
To investigate this issue, we first examine the performance of conservative caching methods and the aggressive caching method quantitatively. As shown in Figure 2, we observe that the caching error from conservative caching methods increases at a relatively slower rate as the caching step grows, while the error in the aggressive caching method accumulates rapidly with each additional caching step, which explains the aforementioned tradeoff. Notably, we find that the caching error of the aggressive and conservative caching exhibit a close value at the first caching step (i.e., timestep 27 in Figure 2) but a significantly larger value in the following caching steps, suggesting that it is possible to have high acceleration ratios and generation quality by leveraging both cache schemes at appropriate caching steps.
Based on this observation, we propose DUal CAching, which leverages the advantages of aggressive and conservative caching methods. In each caching cycle, DuCa first initializes the cache by performing the full computation in the fresh timestep. Then, it performs the one-step aggressive caching for high-ratio acceleration, followed by one-step conservative caching to fix the cache error. Such an alternative caching strategy can be repeated multiple times.
In addition, we further introduce a token-wise caching method as our conservative caching strategy. Token-wise feature caching has been recently proposed to perform computation for only the important tokens in the caching steps. However, the definition of important tokens in the previous work [35] relies on the attention weights in self-attention, making it not compatible with FlashAttention[9] and memory-efficient attention. This problem further improves the memory costs of attention from $O(N)$ to $O(N^{2})$, also resulting in additional latency, making it not applicable. To address this issue, we introduce V-Caching, which approximates the attention weight-based importance score by the norm of the value matrix, providing good compatibility with FlashAttention and memory-efficient Attention while maintaining good generation quality.
Extensive experiments have demonstrated the performance of our method in abundant benchmarks, including text-to-image generation with PixArt-$\alpha$, text-to-video generation with OpenSora, and class-conditioned image generation with DiT. For instance, 1.98×1.98$\times$1.98 × acceleration on PixArt-$\alpha$ and 2.50×2.50$\times$2.50 × acceleration on OpenSora with almost no loss on generation quality.
In summary, our main contributions are as follows:
  1. 1.
    We study the caching error for aggressive feature caching and conservative feature caching, and introduce DuCa, a dual feature caching strategy to perform the two caching methods alternatively, maintaining the generation quality while achieving a higher acceleration ratio.
  2. 2.
    We introduce V-caching as the conservative caching strategy in DuCa, a token-wise feature caching strategy that selects the important tokens via the norm of their value matrix, making it compatible with FlashAttention[9].
  3. 3.
    Extensive experiments have demonstrated DuCa in conditional image generation, text-to-image generation, and text-to-video generation, outperforming previous methods by a large margin. Our codes have been released.
2 Related Works
Transformers in Diffusion Models Diffusion Models (DMs)[16, 30] generate images by gradually adding noise to the image during training and iteratively applying a denoising model to a randomly initialized noise during inference. DMs have recently gained extensive application across various tasks in image generation due to their outstanding performance [27, 1]. Earlier DMs with U-Net-based architectures [16, 27], which also include some Transformer layers, have achieved impressive results. Recently, Diffusion Transformer (DiT) [26] has been constructed by directly stacking multiple DiT Blocks in a cascade, achieving performance in class-conditional image generation that surpasses U-Net-based DMs. Consequently, the DiT architecture has been further extended to several generative fields, yielding exceptional outcomes. For example, the PixArt series [7, 6, 5] and similar text-to-image models incorporate Cross-Attention into the DiT Block to add text conditional information into generation procedure, while the Sora series [4, 18] enhances the DiT Block’s Attention structure to Spatial-Temporal attention, enabling high-quality video generation. However, though success, these DiT-based DMs also entail significantly higher computational costs and time consumption, making deploying such generative models in real-world applications challenging.
Acceleration of Diffusion Models To address the high computational costs commonly associated with DMs including DiT, various acceleration methods have been proposed. Existing acceleration methods can generally be divided into two categories: Designing sampling methods to reduce the number of steps needed to transition from noise to high-quality images, and directly accelerating the denoising model. Sampling-based acceleration methods enhance the efficiency of DMs by simply reducing the sampling steps. DDIM [31] leverages a deterministic, non-Markovian reverse process to significantly reduce generation steps while maintaining high-quality outputs. DPM Solver [22], through analytic derivation, optimizes the reverse process, enabling efficient sampling with fewer steps, while DPM Solver++ [23] further refines this approach by incorporating noise control and diverse sampling techniques to improve quality and adaptability. Rectified Flow [21] employs a recurrent flow framework, utilizing specially designed flow fields to accelerate the diffusion process.
Feature Caching Methods that improve generation efficiency by accelerating the denoising model largely vary depending on the model structure. Acceleration methods for denoising models in different aspects have been proposed and work well, including weight quantization [29], structural pruning [11], knowledge distillation [19] and token merge [3]. Due to the high computational cost of retraining the diffusion models, or even small learnable modules introduce significant costs, limiting model reusability across different scales, resolutions, and tasks, driving interest in training-free acceleration methods. Faster Diffusion [27] improves computational efficiency by caching the output of the encoder module on certain steps, while DeepCache [25] reduces redundant computations by reusing low-resolution feature information in the skip connections of the U-Net structure. Unfortunately, although both are training-free methods, they are specialized feature caching techniques designed specifically for U-Net-based models like Stable Diffusion [27], making them difficult to directly adapt to transformer-based models like DiT [8]. Recently, to address the lack of feature reuse solutions for DiT, $\Delta$-DiT [8] and Pyramid Attention Broadcast (PAB) [33] have been introduced. The former focuses on constructing residuals of DiT outputs across different layers, while the latter emphasizes the importance of various types of attention within DiT Blocks containing multiple attention heads. Learning-to-Cache [24] achieves higher acceleration ratios by employing a learnable router to determine whether computations are needed at each layer, but it incurs prohibitive computational costs. These feature caching methods generally lack attention to finer-grained features. ToCa [35] proposes a method that selects important tokens at both layer and token granularity, allocating more computation to these tokens while caching less important ones, advancing temporal redundancy management to the token level.
The errors introduced by feature caching are primarily attributed to discrepancies in timestep-sensitive information. Comparing the feature caching methods for DiT discussed above, we find that these feature caching approaches for DiT employ different types of feature cache-reuse structures, resulting in significant variations in acceleration ratios while maintaining generation quality. Additionally, they commonly introduce additional computations to mitigate and correct the discrepancies between cached features and the requirements of the current time step. Our proposed method DuCa demonstrates that conservative architectures like ToCa, with their long-range feature reuse capability, can correct timestep information for aggressive methods that struggle with larger timestep intervals. By leveraging the high caching ratios of aggressive methods, DuCa further enhances acceleration efficiency.
3 Method
3.1 Preliminary
Diffusion Models
Diffusion models [16] are designed with two main processes: a forward process, where Gaussian noise is incrementally added to a clean image, and a reverse process, where a standard Gaussian noise is gradually denoised to reconstruct the original image. Letting $t$ represent the timestep and $\beta_{t}$ the noise variance schedule, the conditional probability in the reverse (denoising) process, $p_{\theta}(x_{t-1}\mid x_{t})$, can be modeled as:
$\mathcal{N}\left(x_{t-1};\frac{1}{\sqrt{\alpha_{t}}}\left(x_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(x_{t},t)\right),\beta_{t}\mathbf{I}\right),$ (1)
where $\alpha_{t}=1-\beta_{t}$, $\bar{\alpha}_{t}=\prod_{i=1}^{T}\alpha_{i}$, and $T$ represents the total number of timesteps. Notably, $\epsilon_{\theta}$ is a denoising network parameterized by $\theta$, which takes $x_{t}$ and $t$ as the inputs to predict the noise needed for denoising. In an image generation procedure using $\mathcal{T}$ timesteps, $\epsilon_{\theta}$ must perform inference $\mathcal{T}$ times, which constitutes the majority of computational costs in diffusion models. Recently, numerous studies have shown that implementing $\epsilon_{\theta}$ as a transformer often enhances generation quality.
Figure 3: The overview of DuCa. (a) In the first timestep of a cache cycle, DuCa computes all tokens and stores them in the cache as initialization. Then, DuCa performs a conservative caching step followed by an aggressive caching step, and then another conservative caching step, alternating in this manner to form a complete cache cycle. (b) The proposed V-Caching method selects important tokens for DuCa by first fetching the value matrix for all input tokens that have been previously computed in the attention layers and then computing the norm of the value matrix for each token. The tokens with the largest V-norm values are selected as cached tokens. The computation results for these tokens are directly retrieved from the cache, while the remaining tokens are computed through the neural layer, with their updated results stored back into the cache.
Diffusion Transformers
Diffusion transformers (DiT) are usually composed of stacking groups of self-attention layers $\mathcal{F}_{\text{SA}}$, multilayer perceptron $\mathcal{F}_{\text{MLP}}$, and cross-attention layers $\mathcal{F}_{\text{CA}}$ (for conditional generation). It can be roughly formulated as:
$\mathcal{G}=g_{1}\circ g_{2}\circ\dots\circ g_{L},~{}~{}~{}\text{where}$ (2)
$g_{l}=\mathcal{F}_{\text{SA}}^{l}\circ\mathcal{F}_{\text{CA}}^{l}\circ\mathcal{F}_{\text{MLP}}^{l},$
$\mathcal{G}$, $g_{l}$, and $\mathcal{F}^{l}$ represent the DiT model, the blocks in DiT, and different layers in a DiT block, respectively. $l$ denotes the index of DiT blocks and $L$ denotes the depth of DiT. In diffusion transformers, the input data $\mathbf{x}_{t}$ consists of a sequence of tokens representing different patches in the generated images, expressed as $\mathbf{x}_{t}=\{x_{i}\}_{i=1}^{H\times W}$, where $H$ and $W$ indicate the height and width of the images or their latent code, respectively. Then, the computation between $\mathcal{F}$ and $\mathbf{x}$ can be generally written as $\mathcal{F}(\mathbf{x})=\mathbf{x}+\text{AdaLN}\circ f(\mathbf{x})$, where the first item $\mathbf{x}$ indicates the residual connection and $f$ denotes the MLP or attention layers. AdaLN indicates the adaptive layer normalization [13].
Feature Caching: Feature caching aims to reuse the features that are computed and cached in the past timesteps to skip the computation in future timesteps. It is usually performed in multiple repeated caching cycles. Assuming a caching cycle with $N$ timesteps from the timestep $t$ to $t+N-1$, then feature caching performs the full computation in the timestep $t$ and stores the obtained features in a cache. This progress can be denoted as $\mathcal{C}[l]:=g_{l}(x_{t})$, where $\mathcal{C}$ denotes the cache and “:=” denotes assigning value operation. Then, in timesteps from $t+1$ to $t+N-1$, the computation in these steps is skipped by reusing the cached features, which can be formulated as $g_{l}(x_{t+i}):=\mathcal{C}[l],i\leq N-1$. For simplicity, in this paper, the timestep $t$ and the timesteps in $[t+1,t+N-1]$ are named the freshing step and the caching steps, respectively.
3.2 Dual Feature Caching
The proposed dual feature caching aims to perform aggressive caching and conservative caching alternatively in the caching steps. Hence, in this section, we first introduce the details of the two caching strategies.
Aggressive Caching in DuCa
As introduced in Figure 1(b), aggressive caching directly replaces the output of the DiT block with the cached feature which is computed in previous layers, which means that all computations of the first $l$ layers can be skipped. This can be expressed as $g_{l}:=\mathcal{C}[l]$, where $\mathcal{C}[l]$ denotes the cached feature at the $l_{th}$ layer. Then, the computation with aggressive caching can be written as:
$\mathcal{G}(x)_{t}:=g_{l+1}\circ...\circ g_{L}^{t}(\mathcal{C}[l]).$ (3)
To achieve the maximal acceleration with aggressive caching, we set $l=L-1$ in most of our experiments, which indicates almost skipping all the layers in the caching step.
Conservative Caching in DuCa
As shown in Figure 1(a), conservative caching does not cache and reuse all the features. Instead, it caches and reuses only the features without the residual connection, which allows it to allocate the computation to some layers in the caching steps to correct the cached features. Recently, ToCa [35] introduced a token-wise approach that selectively reuses features based on the importance of each token and its sensitivity to caching, which enables us to perform computation only in the important tokens. Following ToCa, we first calculate an importance score $\mathcal{S}$ for each token $x_{i}$ and then use this score to divide the index set of tokens $\mathcal{I}$ into two complementary sets: $\mathcal{I}_{\text{Cache}}$, for tokens to be cached, and $\mathcal{I}_{\text{Compute}}$, for tokens to be computed. Subsequently, the computation for the $i_{th}$ token $x_{i}$ on the $l_{th}$ layer can be expressed as:
$\gamma_{i}f^{l}(x_{i})+(1-\gamma_{i})\mathcal{C}[l](x_{i}),$ (4)
where $\gamma_{i}=0$ for $i\in\mathcal{I}_{\text{Cache}}$ and $\gamma_{i}=1$ for $i\in\mathcal{I}_{\text{Compute}}$. Here, $\mathcal{C}[l](x_{i})$ denotes retrieving the cached value of $x_{i}$ from the conservative cache $\mathcal{C}$ at layer $l$, which incurs no computation cost, enabling partial omission of calculations. After this selective computation in the token dimension, the computed features are updated to $\mathcal{C}[l]$ for future reusing.
Value-based Token-wise Feature Caching:
As the previous token-wise conservative caching method, ToCa offers two attention map-based selection strategies that have shown good effectiveness. However, the attention map-based selection strategy is incompatible with another commonly used attention-optimization method FlashAttention[9] and memory-efficient attention, which can not provide the attention scores but are capable of reducing the memory costs of attention from $O(N^{2})$ to $O(N)$ and significant accelerating the attention layers. This conflict between attention scores-based token selection with this optimized attention computation makes ToCa less practical in application and raises the requirements for attention score-free token selection methods.
Figure 4 shows the relation between the norm of the value matrix and the attention scores of each token, which demonstrates that tokens with high attention scores in the self-attention map tend to have relatively smaller norms in their value matrix. This suggests that we can directly select tokens with smaller norms in their value matrix as substitutes for those with high attention map scores. Concretely, the value matrix of each token is produced by a linear layer in attention layers as follows: $v_{i}=\text{MLP}_{v}(x_{i})$, where $v_{i}$ is value vector for the $i^{th}$ token. Therefore, V-Caching gives the selection rule for $\mathcal{I}_{\text{Cache}}$ as follows:
$\mathcal{I}_{\text{Cache}}=\mathop{\arg\max}_{\{i_{1},i_{2},\dots,i_{n}\}\subseteq\{1,2,\dots,N\}}\{\|v_{i_{1}}\|_{2},\|v_{i_{2}}\|_{2},\cdots,\|v_{i_{n}}\|_{2}\},$ (5)
where the $n$ is the total number of selected tokens, defined by $n=\lfloor R\%\times N\rfloor$.
3.3 Overall Framework
Built on the deeper understanding of aggressive and conservative caching, DuCa introduces a caching scheme that alternates conservative and aggressive caching in different steps. V-Caching serves as the conservative caching strategy, as shown in Figure 3. This design leverages the realignment capability of token-wise conservative caching to correct the misaligned timestep information occurring in aggressive caching steps. Therefore, it enables the high local compression rate of aggressive caching to be applied to more steps while further reducing redundant computations in conservative caching steps. In practice, we arrange token-wise conservative caching steps on odd-numbered steps following a fresh step, and aggressive caching steps on even-numbered steps.
Figure 4: The relation between the self-attention score and the norm of the value matrix for each token. Tokens with the largest attention score tend to have a lower norm in the value matrix.
4 Experiments
4.1 Experiment Settings
4.1.1 Model Configurations
Experiments are conducted on three widely known DiT-based models on the three generation tasks with the correlating default samplers, including PixArt-$\alpha$ [7] with 20 DPM Solver++ [23] steps for text-to-image generation, OpenSora [34] with 30 reflected flow steps [21] for text-to-video generation, and DiT-XL/2 [26] with 50 DDIM [31] steps for class-conditional image generation. All the experiments are conducted with NVIDIA H800 80GB GPUs. See More details in our appendix.
4.1.2 Evaluation and Metrics
Text-to-image generation We randomly select 30,000 captions from COCO-2017 datasets [20] to generate the 30,000 $256\times 256$ images. FID-30k is computed using the 30k generated images and the real images from COCO-2017 to evaluate the quality of text-to-image generation. Besides, the CLIP score [14] is computed between the 30k prompt-image pairs to evaluate the text-image alignment. All the CLIP scores are tested using the Vit-large-14 model.
Text-to-video generation VBench [17] evaluation framework is applied to evaluate the performance of the text-to-video model. With 5 videos per prompt and 950 prompts from the framework, 4,750 videos are generated to evaluate the 16 aspects proposed by VBench. The generated videos are 480p, 9:16, and 2s videos.
Class-conditional image generation 50,000 $256\times 256$ images uniformly sampled from the 1,000 classes in ImageNet[10] are generated to evaluate the FID-50k [15]. Besides, sFID, Precision, and Recall are introduced as supplementary metrics.
Table 1: Quantitative comparison of text-to-image generation on MS-COC02017 with PixArt-$\alpha$ and 20 DPM++ sampling steps.
[…]
PixArt-$\alpha$ [7] 0.242 11.88 1.00$\times$ 28.12 16.29
$50\%$ steps 0.121 5.94 2.00$\times$ 37.57 15.83
FORA [28] 0.178 5.66 1.98$\times$ 29.67 16.40
ToCa[35] 0.167 6.05 1.96$\times$ 28.38 16.43
DuCa(C-A score) 0.157 5.99 1.98$\times$ 27.98 16.44
DuCa(V-Caching) 0.088 5.99 1.98$\times$ 28.05 16.42
Table 2: Quantitative comparison in text-to-video generation on VBench. *Results are from PAB [33]. PAB1-3 indicate PAB with different hyper-parameters.
[…]
OpenSora [34] 54.83 3283.20 1.00$\times$ 79.41
[…]
PAB1∗ [33] 41.05 2657.70 1.24$\times$ 78.51
PAB2∗ [33] 39.96 2615.15 1.26$\times$ 77.64
PAB3∗ [33] 38.26 2558.25 1.28$\times$ 76.95
$50\%$ steps 28.85 1641.60 2.00$\times$ 76.78
FORA [28] 33.19 1751.32 1.87$\times$ 76.91
ToCa[35] 29.03 1394.03 2.36$\times$ 78.59
DuCa(C-A score) 26.77 1315.62 2.50$\times$ 78.63
DuCa(V-Caching) 19.16 1315.62 2.50$\times$ 78.54
Table 3: Quantitative comparison on class-to-image generation on ImageNet with DiT-XL/2.
Method Latency(s) $\downarrow$ FLOPs(T) $\downarrow$ Speed $\uparrow$ sFID $\downarrow$ FID $\downarrow$ Precision $\uparrow$ Recall $\uparrow$
DiT-XL/2-G (cfg $=1.50$) 2.012 118.68 1.00$\times$ 4.98 2.31 0.82 0.58
DDIM-50 steps 0.417 23.74 1.00$\times$ 4.40 2.43 0.80 0.59
DDIM-25 steps 0.209 11.87 2.00$\times$ 4.74 3.18 0.79 0.58
DDIM-20 steps 0.167 9.49 2.50$\times$ 5.15 3.81 0.78 0.58
DDIM-17 steps 0.142 8.07 2.94$\times$ 5.76 4.58 0.77 0.56
FORA [28] 0.197 4.29 2.77$\times$ 6.43 3.88 0.79 0.54
ToCa(a) 0.195 8.73 2.72$\times$ 5.14 3.64 0.79 0.55
ToCa(b) 0.214 10.23 2.32$\times$ 4.70 3.05 0.79 0.57
DuCa(a)(S-A score) 0.184 8.76 2.71$\times$ 4.66 3.19 0.79 0.57
DuCa(b)(S-A score) 0.200 9.58 2.48$\times$ 4.66 3.05 0.80 0.57
DuCa(a)(V-Caching) 0.177 8.76 2.71$\times$ 4.68 3.19 0.80 0.57
DuCa(b)(V-Caching) 0.198 9.58 2.48$\times$ 4.69 3.06 0.80 0.57
4.2 Quantitative and Qualitative Results
Figure 5: Visualization results for different acceleration methods on PixArt-$\alpha$.
Figure 6: Visualization results for different methods on OpenSora (videos available in the supplementary material).
4.2.1 Results on text-to-image generation
As shown in Table 1, we compared the DuCa under two different token selection methods: Cross-Attention score from ToCa and the proposed V-norm score with three training-free acceleration methods: ToCa[35], the 10 DPM++ sampling steps and FORA [28], a layer-wise conservative caching method.
In terms of generation quality, the quantitative results demonstrate that under an acceleration ratio close to 2.0, the DuCa methods achieve the lowest FID-30k score on the MS-COCO2017 generation task, with almost no loss in generation quality compared to the original non-accelerated approach. For instance, the DuCa method using the Cross-Attention score reduces the FID-30k by 0.40 compared to the ToCa method with the same acceleration ratio, while the DuCa method using the V-norm score also decreases the FID-30k by 0.33, demonstrating that our alternating acceleration strategy effectively enhances generation quality.
In terms of evaluating the text-image alignment metric, the performance of the DuCa method is very close to that of ToCa and generally superior to the approach of directly setting $50\%$ generation steps. For example, the DuCa method with Cross-Attention score and V-norm score achieves CLIP scores on MS-COCO2017 that are 0.61 and 0.59 higher, respectively, compared to the $50\%$ generation steps method with the same acceleration ratio. These scores are even slightly better than those of the original non-accelerated PixArt-$\alpha$ model, with improvements of 0.15 and 0.13, respectively, demonstrating better controllability.
In addition, thanks to the acceleration brought by the compatibility with FlashAttention [9], the latency of the DuCa method with V-norm score is reduced by 0.069 compared to the DuCa method with Attention score, representing an approximate $43.9\%$ reduction, which is a significant improvement in acceleration.
4.2.2 Results on text-to-video generation
Figure 7: VBench metrics and acceleration ratio of proposed DuCa and other methods
The generation results of $\Delta$-DiT, T-GATE[32], PAB[33], 15 rflow sampling steps, ToCa[35], FORA [28] and proposed DuCa on the text-to-video model OpenSora[34] are shown in Table 2.
As shown in Table 2, DuCa achieves a further acceleration from 2.36 $\times$ (as with ToCa) to 2.50 $\times$ with virtually no loss in generation quality. Additionally, despite having the lowest computational cost among all the compared methods, DuCa still achieves the highest VBench score. This demonstrates that the proposed strategy in DuCa, which alternates between aggressive and conservative caching steps, effectively eliminates the redundant components present in a fully conservative caching approach.
Additionally, comparing the generation results of the Cross-Attention and V-Caching token selection strategies, we only observe a slight decrease of 0.09 in the VBench score with the V-Caching strategy. More significantly, V-Caching reduces generation time by 7.61 seconds, approximately $28.4\%$. This indicates that DuCa effectively leverages V-Caching to make the acceleration method compatible with FlashAttention[9], achieving substantial reductions in generation time. In practice, using V-Caching will be more advantageous for rapid inference.
In Figure 7, we further present the performance of DuCa and the aforementioned methods across various dimensions of the VBench metric. DuCa shows results consistent with the non-accelerated original OpenSora model across multiple metrics, including but not limited to Overall Consistency, Temporal Style, Scene, and Object Class. Additionally, the Cross-Attention and V-Caching strategies introduce only a slight difference in Spatial Relationship, further underscoring the exceptional performance of DuCa and the proposed V-Caching in practical scenarios.
4.2.3 Results on class-conditional image generation
In Table 3, we compare ToCa[35], FORA [28], the direct reduction of corresponding DDIM steps, and the proposed DuCa under Self-Attention score and V-Caching token selection strategies. The experimental results indicate that with comparable FID scores, representing generation quality, DuCa further enhances acceleration. For instance, DuCa(b) with Self-Attention score achieves a similar FID-50k as ToCa(b), but the acceleration ratio increases from 2.36 to 2.48. Additionally, under a higher acceleration scenario with an acceleration ratio around 2.7, DuCa achieves an FID-50k of 3.19, further reducing the FID-50k by 0.45 compared to ToCa’s 3.64. This demonstrates that DuCa effectively reallocates computation to more critical areas.
4.3 Ablation Studies
Table 4: Ablation Study for different token selection methods for DuCa on PixArt-$\alpha$.
Methods MS-COCO2017 Latency(s) $\downarrow$ FLOPs(T) $\downarrow$
FID-30k $\downarrow$ CLIP $\uparrow$
Random 28.34 16.42 0.089 5.99
Cross-Attention 27.98 16.44 0.157 5.99
V-Caching 28.05 16.42 0.088 5.99
Table 5: Comparison on different caching methods on PixArt-$\alpha$.
Methods MS-COCO2017 FLOPs(T) $\downarrow$
FID-30k $\downarrow$ CLIP $\uparrow$
Conservative-only 27.92 16.39 7.84
Aggressive-only 30.51 16.39 3.59
DuCa 28.05 16.42 5.99
As shown in Table 4, we compare Random, Cross-Attention based, and V-Caching token selection methods on PixArt-$\alpha$ in terms of latency, FLOPs, FID-30k, and CLIP score. It can be observed that the cross-Attention-based selection method achieves the best performance in generation quality and text-image alignment, but due to its incompatibility with FlashAttention[9], its latency is as high as 0.157 seconds—nearly twice that of V-Caching and Random selection. In contrast, V-Caching achieves nearly the same, only slightly lower, FID-30k and CLIP scores, along with faster generation speed. Despite the minimal impact of the three token selection methods on computational load, the FlashAttention[9]-compatible V-Caching method shows significantly higher acceleration due to computational optimizations. Thus, V-Caching stands out as a superior generation scheme that effectively balances quality and efficiency.
As shown in Table 5, we compare the Conservative-only, Aggressive-only, and alternating DuCa structures on PixArt-$\alpha$ in terms of FLOPs, FID-30k, and CLIP score. The Conservative-only structure appears to have the best generation quality but incurs an excessively high computational cost. In contrast, Aggressive-only reduces computation to nearly half of Conservative-only but suffers significantly in generation quality. Our proposed DuCa structure achieves a computational load between the two, with only a slight 0.13 increase in FID-30k compared to the high-computation Conservative-only approach, while improving text-image alignment. This indicates that DuCa effectively eliminates most redundant computation present in Conservative-only.
5 Conclusion
This paper first gives an in-depth study of the caching error on aggressive and conservative methods, and then introduces dual caching DuCa, which performs aggressive and conservative alternatively to leverage their advantages in different caching steps. Besides, V-Caching is introduced to perform token-wise conservative caching based on the norm of the value matrix of the tokens, making it compatible with default optimized attention computation such as FlashAttention, allowing feature caching to be practical in real applications. We hope this paper can attract more attention to feature caching for its real benefits, to make it simple but effective, instead of being complex.
Supplementary Material
6 Experimental Details
In this section, more details of the experiments are provided.
6.1 Model Configuration
As mentioned in 4.1.1 , experiments on three models from different tasks, PixArt-$\alpha$[7] for text-to-image generation, OpenSora[4] for text-to-video generation, and DiT[26] for class-conditional image generation, are presented. In this section, a more detailed hyperparameter configuration scheme is provided.
  • PixArt-$\alpha$: The average cache cycle length is set to $\mathcal{N}=3$ with an average caching ratio of $R=70\%$ for ToCa[35], and $\mathcal{N}=3$ with $\mathcal{R}=25\%$ for the DuCa method to maintain a similar acceleration ratio. Additionally, in the FORA [28] scheme, $\mathcal{N}=2$. Conservative caching steps are configured to occur on odd-numbered steps following fresh steps, while aggressive caching steps occur on even-numbered steps following fresh steps. The proportion of conservative caching steps across layers and steps is identical to that in ToCa[35].
  • OpenSora: The average cache cycle length is $\mathcal{N}=3$ for temporal attention, spatial attention, and MLP, and $\mathcal{N}=6$ for cross-attention, with an average caching ratio $\mathcal{R}=85\%$ exclusively for the MLP in ToCa[35] and DuCa, while all other modules use a caching ratio of $100\%$. In the FORA [28]scheme, all modules are set to $\mathcal{N}=2$.
    For $\textbf{PAB}^{1*}$ [33], temporal attention is set to $\mathcal{N}=2$, spatial attention to $\mathcal{N}=4$, cross-attention to $\mathcal{N}=6$, and the MLP module is fully computed. For $\textbf{PAB}^{2*}$ [33], $\mathcal{N}$ is set to 3, 5, and 7 for temporal, spatial, and cross-attention, respectively. For $\textbf{PAB}^{3*}$ [33], $\mathcal{N}$ is set to 5, 7, and 9 for temporal, spatial, and cross-attention, respectively.
    For DuCa, conservative caching steps are set on odd-numbered steps following fresh steps, while aggressive caching steps are set on even-numbered steps following fresh steps. The proportion of conservative caching steps across layers and steps is identical to that in ToCa[35].
  • DiT: The average cache cycle length is set to $\mathcal{N}=4$ with an average caching ratio of $R=93\%$ for ToCa[35](a), $\mathcal{N}=3$ with $R=93\%$ for ToCa[35](b), and $\mathcal{N}=3$ with $R=95\%$ for both DuCa(a) and (b). Additionally, the average cache cycle length for FORA [28]is set to $\mathcal{N}=3$.
    Furthermore, for DuCa(a), conservative caching steps are set on even-numbered steps following fresh steps, while aggressive caching steps are set on odd-numbered steps. In contrast, for DuCa(b), the arrangement is reversed.
    Since the average cache cycle length $\mathcal{N}=3$ here, DuCa(a), which places aggressive caching steps immediately after the fresh step, achieves a higher acceleration ratio compared to DuCa(b), which places conservative caching steps in the first position after the fresh step.
6.2 Anonymous Page for video presentation
To further demonstrate the superiority of DuCa in video generation, we have created an anonymous GitHub page. Please visit https://duca2024.github.io/DuCa/ for a more detailed view. Besides, the videos are also available through the Supplementary Material.

### v2

@@
+Rethinking Token-wise Feature Caching: Accelerating Diffusion Transformers with Dual Feature Caching
+Rethinking Token-wise Feature Caching:
 Accelerating Diffusion Transformers with Dual Feature Caching
-Abstract
+Chang Zou1,*, Shikang Zheng1,2,*, Evelyn Zhang1, Runlin Guo3, Haohang Xu4, Zhengyi Shi1
+Conghui He5
+, Xuming Hu6, Linfeng Zhang1,†
+1
+Shanghai Jiao Tong University 2 South China University of Technology 3 Beihang University
+4 Huawei Technologies Ltd 5 Shanghai AI Lab 6 Hong Kong University of Science and Technology (Guangzhou)
+* Equal contribution.
+Denotes the corresponding author.
+I Abstract
 Diffusion Transformers (DiT) have become the dominant methods in image and video generation yet still suffer substantial computational costs.
 As an effective approach for DiT acceleration, feature caching methods are designed to cache the features of DiT in previous timesteps and reuse them in the next timesteps, allowing us to skip the computation in the next timesteps.
-However, on the one hand, aggressively reusing all the features cached in previous timesteps leads to a severe drop in generation quality.
-On the other hand, conservatively caching only the features in the redundant layers or tokens but still computing the important ones successfully preserves the generation quality but results in reductions in acceleration ratios.
-Observing such a tradeoff between generation quality and acceleration performance, this paper begins by quantitatively studying the accumulated error from cached features.
-Surprisingly, we find that aggressive caching does not introduce significantly more caching errors in the caching step, and the conservative feature caching can fix the error introduced by aggressive caching.
-Thereby, we propose a dual caching strategy that adopts aggressive and conservative caching iteratively, leading to significant acceleration and high generation quality at the same time.
-Besides, we further introduce a V-caching strategy for token-wise conservative caching, which is compatible with flash attention and requires no training and calibration data.
-Our codes have been released in Github:
-Code: https://github.com/Shenyi-Z/DuCa
-1 Introduction
-Diffusion Models (DMs)[16] have shown remarkable success in both image [27] and video generation [2].
-Traditional DMs often use U-Net denoising structures, such as Stable Diffusion (SD)[27] and Stable Video Diffusion (SVD)[2], which have already demonstrated impressive results.
-Most recently, with the introduction of Diffusion Transformers (DiT)[26], the quality of visual generation has further significantly improved, propelling the development of numerous downstream applications [12].
+Among them, token-wise feature caching has been introduced to perform different caching ratios for different tokens in DiTs, aiming to skip the computation for unimportant tokens while still computing the important ones.
+In this paper, we propose to carefully check the effectiveness in token-wise feature caching with the following two questions: (1) Is it really necessary to compute the so-called ”important” tokens in each step?
+(2) Are so-called important tokens really important?
+Surprisingly, this paper gives some counter-intuition answers, demonstrating that consistently computing the selected “important tokens” in all steps is not necessary.
+The selection of the so-called “important tokens” is often ineffective, and even sometimes shows inferior performance than random selection.
+Based on these observations, this paper introduces dual feature caching referred to as DuCa, which performs aggressive caching strategy and conservative caching strategy iteratively and selects the tokens for computing randomly.
+Extensive experimental results demonstrate the effectiveness of our method in DiT, PixArt, FLUX, and OpenSora, demonstrating significant improvements than the previous token-wise feature caching.
+II Introduction
+Diffusion Models (DMs)[ho2020DDPM] have shown remarkable success in both image [rombach2022SD] and video generation [blattmann2023SVD].
+Traditional DMs often use U-Net denoising structures, such as Stable Diffusion (SD)[rombach2022SD] and Stable Video Diffusion (SVD)[blattmann2023SVD], which have already demonstrated impressive results.
+Most recently, with the introduction of Diffusion Transformers (DiT)[peebles2023dit], the quality of visual generation has significantly improved further, propelling the development of numerous downstream applications [Feng2024DiT4EditDT].
 However, the quality improvement is also accompanied by a significant increase in computational demands, reducing the inference efficiency of transformers and making their deployment in practical scenarios a pressing challenge.
 Recently, feature caching has become one of the most popular techniques for diffusion acceleration.
-Motivated by the high similarity between features in the adjacent timesteps, aggressive feature caching caches the features computed in a freshing timestep, and then reuses them in the following several caching steps, as shown in Figure 1(b).
+Motivated by the high similarity between features in the adjacent timesteps, aggressive feature caching caches the features computed in a freshing timestep, and then reuse them in the following several caching steps, as shown in Figure 1(b).
 Such a fresh-then-cache strategy enables the diffusion model to save computational costs in the caching step and thus leads to significant acceleration.
 However, since the similarity of features in non-adjacent timesteps decreases quickly, the error introduced by aggressive feature caching can be accumulated at an exponential speed, leading to a significant drop in generation quality, and making aggressive feature caching impractical.
-Figure 1: Comparison between conservative caching (a) and aggressive caching (b): Conservative caching skips fewer layers than aggressive caching to maintain the generation quality, but also leads to lower acceleration ratios.
-$t$ and $t-1$ denote the refreshing and caching timesteps, respectively.
-To solve this problem, recent works [8, 24, 28, 35] introduce a few computations during the caching steps in certain layers and tokens, which make the cached features in previous steps be corrected for the current caching step.
-These conservative feature caching methods reduce the accumulated error from feature caching but also make it difficult to achieve a higher acceleration ratio.
-Such a tradeoff between acceleration ratio and generation quality becomes the most essential problem for feature caching methods.
+Figure 1: Comparison between conservative caching (a) and aggressive caching (b): Conservative caching skips some tokens instead of all tokens in each layer to maintain the generation quality, but also leads to lower acceleration ratios.
+$t$ and t1t-1 denote the refreshing and caching timesteps, respectively.
+To solve this problem, recent work [chen2024delta-dit, ma2024l2c, selvaraju2024fora, zou2024accelerating] introduces some computations during the caching steps in certain layers and tokens, which makes the cached features in previous steps corrected for the current caching step.
+Among them, token-wise feature caching (ToCa) has been proposed to selectively compute the important tokens while skipping the unimportant tokens with the cache, reducing the accumulated error from feature caching, but also making it difficult to achieve a higher acceleration ratio.
+In this paper, we rethink and distill ToCa.
+Specifically, we focus on the following two questions.
+(I) Is it necessary to compute the important tokens in all caching steps?
+ToCa computes the important tokens in each caching step to reduce the accumulated error from feature caching.
+Figure 2 shows the caching error from the aggressive (marked in green) and conservative caching (ToCa, marked in blue).
+It is observed that the caching error from ToCa increases slowly as the caching step grows, while the error in the aggressive caching method accumulates rapidly with each caching step, which verifies the effectiveness of ToCa.
+However, we find that the caching error of the aggressive and conservative caching exhibit a close value at the first caching step (i.e., timestep 27 in Figure 2), suggesting that computing the important tokens in ToCa is not always necessary.
+Concretely, when the error from caching has not been accumulated yet, computing important tokens does not bring benefits and can be removed for better efficiency.
+Solution: Dual Feature Caching: Based on this observation, we propose DUal CAching, which leverages the advantages of aggressive and conservative caching methods.
+In each caching cycle, DuCa initializes the cache by performing the full computation in the fresh timestep.
+Then, it performs the one-step aggressive caching for high-ratio acceleration, followed by one-step conservative caching to fix the cache error.
+This alternative caching strategy can be repeated multiple times to achieve the best balance between efficiency and quality.
+(II) Are so-called important tokens really important?
+The basic assumption of ToCa is that there exist some tokens that are sensitive or robust to feature caching.
+Based on this motivation, ToCa introduces four kinds of scores to describe the importance of each token based on their attention scores, spatial position, and frequency of being cached, respectively.
+Unfortunately, these scores not only introduce additional computation costs, but also lead to the incompatibility of FlashAttention, one of the default efficient operators in hardware.
+Hence, it results in around 100% increments in latency and improves the memory overhead from $O(N)$ to $O(N^{2})$.
+To solve this problem, we carefully study the effectiveness of each score in ToCa.
+Surprisingly, as shown in Table I and will be discussed later, we find that selecting tokens with the highest attention scores in ToCa sometimes shows inferior performance than random token selection, indicating previous methods may misunderstand the importance of tokens.
+Solution: Selecting tokens with low similarity: After carefully trying all kinds of importance scores for token selection, we find that choosing a subset of tokens that has the smallest similarity to other tokens is the only solution to beat random token selection.
+On the other hand, choosing a subset of tokens with maximal similarity to each other almost achieves the worst performance that we can have.
+Based on this observation, we argue that token selection should focus on their duplication instead of their importance.
+Thus, random selection can be a good choice since randomness guarantees choosing tokens with different semantic information and leads to no additional computation or memory overhead.
 Figure 2: The caching error and acceleration ratio of three feature caching strategies.
 The caching error is defined as the $L2$-norm distance between $x_{t}$ computed with and without feature caching.
-To investigate this issue, we first examine the performance of conservative caching methods and the aggressive caching method quantitatively.
-As shown in Figure 2, we observe that the caching error from conservative caching methods increases at a relatively slower rate as the caching step grows, while the error in the aggressive caching method accumulates rapidly with each additional caching step, which explains the aforementioned tradeoff.
-Notably, we find that the caching error of the aggressive and conservative caching exhibit a close value at the first caching step (i.e., timestep 27 in Figure 2) but a significantly larger value in the following caching steps, suggesting that it is possible to have high acceleration ratios and generation quality by leveraging both cache schemes at appropriate caching steps.
-Based on this observation, we propose DUal CAching, which leverages the advantages of aggressive and conservative caching methods.
-In each caching cycle, DuCa first initializes the cache by performing the full computation in the fresh timestep.
-Then, it performs the one-step aggressive caching for high-ratio acceleration, followed by one-step conservative caching to fix the cache error.
-Such an alternative caching strategy can be repeated multiple times.
-In addition, we further introduce a token-wise caching method as our conservative caching strategy.
-Token-wise feature caching has been recently proposed to perform computation for only the important tokens in the caching steps.
-However, the definition of important tokens in the previous work [35] relies on the attention weights in self-attention, making it not compatible with FlashAttention[9] and memory-efficient attention.
-This problem further improves the memory costs of attention from $O(N)$ to $O(N^{2})$, also resulting in additional latency, making it not applicable.
-To address this issue, we introduce V-Caching, which approximates the attention weight-based importance score by the norm of the value matrix, providing good compatibility with FlashAttention and memory-efficient Attention while maintaining good generation quality.
-Extensive experiments have demonstrated the performance of our method in abundant benchmarks, including text-to-image generation with PixArt-$\alpha$, text-to-video generation with OpenSora, and class-conditioned image generation with DiT.
-For instance, 1.98×1.98$\times$1.98 × acceleration on PixArt-$\alpha$ and 2.50×2.50$\times$2.50 × acceleration on OpenSora with almost no loss on generation quality.
+Extensive experiments have demonstrated the performance of our method in abundant benchmarks, including text-to-image generation with FLUX.1-dev, PixArt-$\alpha$, text-to-video generation with OpenSora, and class-conditioned image generation with DiT.
+For instance, $3.45\times$ acceleration on FLUX.1-dev and $2.50\times$ acceleration on OpenSora with almost lossless quality.
 In summary, our main contributions are as follows:
   1. 1.
-    We study the caching error for aggressive feature caching and conservative feature caching, and introduce DuCa, a dual feature caching strategy to perform the two caching methods alternatively, maintaining the generation quality while achieving a higher acceleration ratio.
+    We rethink the necessity of computing “important” tokens in token-wise feature caching, and introduce DuCa, a dual feature caching strategy to perform the two caching methods alternatively, maintaining the generation quality while achieving a higher acceleration ratio.
   2. 2.
-    We introduce V-caching as the conservative caching strategy in DuCa, a token-wise feature caching strategy that selects the important tokens via the norm of their value matrix, making it compatible with FlashAttention[9].
+    We rethink the tokens selection methods in token-wise feature caching, finding that previous selection methods show inferior performance than random token selection, and explaining this phenomenon from the perspective of duplication.
+By just replacing previous methods with random selection, we achieve good compatibility with FlashAttention, leading to $\sim 2\times$ speed up in latency.
   3. 3.
@@
 Our codes have been released.
-2 Related Works
-Transformers in Diffusion Models Diffusion Models (DMs)[16, 30] generate images by gradually adding noise to the image during training and iteratively applying a denoising model to a randomly initialized noise during inference.
-DMs have recently gained extensive application across various tasks in image generation due to their outstanding performance [27, 1].
-Earlier DMs with U-Net-based architectures [16, 27], which also include some Transformer layers, have achieved impressive results.
-Recently, Diffusion Transformer (DiT) [26] has been constructed by directly stacking multiple DiT Blocks in a cascade, achieving performance in class-conditional image generation that surpasses U-Net-based DMs.
-Consequently, the DiT architecture has been further extended to several generative fields, yielding exceptional outcomes.
-For example, the PixArt series [7, 6, 5] and similar text-to-image models incorporate Cross-Attention into the DiT Block to add text conditional information into generation procedure, while the Sora series [4, 18] enhances the DiT Block’s Attention structure to Spatial-Temporal attention, enabling high-quality video generation.
-However, though success, these DiT-based DMs also entail significantly higher computational costs and time consumption, making deploying such generative models in real-world applications challenging.
+III Related Works
 Acceleration of Diffusion Models To address the high computational costs commonly associated with DMs including DiT, various acceleration methods have been proposed.
@@
 Sampling-based acceleration methods enhance the efficiency of DMs by simply reducing the sampling steps.
-DDIM [31] leverages a deterministic, non-Markovian reverse process to significantly reduce generation steps while maintaining high-quality outputs.
-DPM Solver [22], through analytic derivation, optimizes the reverse process, enabling efficient sampling with fewer steps, while DPM Solver++ [23] further refines this approach by incorporating noise control and diverse sampling techniques to improve quality and adaptability.
-Rectified Flow [21] employs a recurrent flow framework, utilizing specially designed flow fields to accelerate the diffusion process.
-Feature Caching Methods that improve generation efficiency by accelerating the denoising model largely vary depending on the model structure.
-Acceleration methods for denoising models in different aspects have been proposed and work well, including weight quantization [29], structural pruning [11], knowledge distillation [19] and token merge [3].
-Due to the high computational cost of retraining the diffusion models, or even small learnable modules introduce significant costs, limiting model reusability across different scales, resolutions, and tasks, driving interest in training-free acceleration methods.
-Faster Diffusion [27] improves computational efficiency by caching the output of the encoder module on certain steps, while DeepCache [25] reduces redundant computations by reusing low-resolution feature information in the skip connections of the U-Net structure.
-Unfortunately, although both are training-free methods, they are specialized feature caching techniques designed specifically for U-Net-based models like Stable Diffusion [27], making them difficult to directly adapt to transformer-based models like DiT [8].
-Recently, to address the lack of feature reuse solutions for DiT, $\Delta$-DiT [8] and Pyramid Attention Broadcast (PAB) [33] have been introduced.
+DDIM [songDDIM] leverages a deterministic, non-Markovian reverse process to significantly reduce generation steps while maintaining high-quality outputs.
+DPM Solver [lu2022dpm], through analytic derivation, optimizes the reverse process, enabling efficient sampling with fewer steps, while DPM Solver++ [lu2022dpm++] further refines this approach by incorporating noise control and diverse sampling techniques to improve quality and adaptability.
+Rectified Flow [refitiedflow] employs a recurrent flow framework, utilizing specially designed flow fields to accelerate the diffusion process.
+Feature Caching Due to the high computational cost of retraining the diffusion models, or even small learnable modules introduce significant costs, limiting model reusability across different scales, resolutions, and tasks, driving interest in training-free acceleration methods.
+Faster Diffusion [rombach2022SD] improves computational efficiency by caching the output of the encoder module on certain steps, while DeepCache [ma2024deepcache] reduces redundant computations by reusing low-resolution feature information in the skip connections of the U-Net structure.
+Unfortunately, although both are training-free methods, they are specialized feature caching techniques designed specifically for U-Net-based models like Stable Diffusion [rombach2022SD], making them difficult to directly adapt to transformer-based models like DiT [chen2024delta-dit].
+Recently, to address the lack of feature reuse solutions for DiT, $\Delta$-DiT [chen2024delta-dit] and Pyramid Attention Broadcast (PAB) [zhao2024PAB] have been introduced.
 The former focuses on constructing residuals of DiT outputs across different layers, while the latter emphasizes the importance of various types of attention within DiT Blocks containing multiple attention heads.
-Learning-to-Cache [24] achieves higher acceleration ratios by employing a learnable router to determine whether computations are needed at each layer, but it incurs prohibitive computational costs.
+Learning-to-Cache [ma2024l2c] achieves higher acceleration ratios by employing a learnable router to determine whether computations are needed at each layer, but it incurs prohibitive computational costs.
 These feature caching methods generally lack attention to finer-grained features.
-ToCa [35] proposes a method that selects important tokens at both layer and token granularity, allocating more computation to these tokens while caching less important ones, advancing temporal redundancy management to the token level.
-The errors introduced by feature caching are primarily attributed to discrepancies in timestep-sensitive information.
-Comparing the feature caching methods for DiT discussed above, we find that these feature caching approaches for DiT employ different types of feature cache-reuse structures, resulting in significant variations in acceleration ratios while maintaining generation quality.
-Additionally, they commonly introduce additional computations to mitigate and correct the discrepancies between cached features and the requirements of the current time step.
-Our proposed method DuCa demonstrates that conservative architectures like ToCa, with their long-range feature reuse capability, can correct timestep information for aggressive methods that struggle with larger timestep intervals.
-By leveraging the high caching ratios of aggressive methods, DuCa further enhances acceleration efficiency.
-3 Method
-3.1 Preliminary
+ToCa [zou2024accelerating] proposes a method that selects important tokens at both layer and token granularity, allocating more computation to these tokens while caching less important ones, advancing temporal redundancy management to the token level.
+IV Method
+IV-A Preliminary
 Diffusion Models
-Diffusion models [16] are designed with two main processes: a forward process, where Gaussian noise is incrementally added to a clean image, and a reverse process, where a standard Gaussian noise is gradually denoised to reconstruct the original image.
+Diffusion models [ho2020DDPM] are designed with two main processes: a forward process, where Gaussian noise is incrementally added to a clean image, and a reverse process, where a standard Gaussian noise is gradually denoised to reconstruct the original image.
 Letting $t$ represent the timestep and $\beta_{t}$ the noise variance schedule, the conditional probability in the reverse (denoising) process, $p_{\theta}(x_{t-1}\mid x_{t})$, can be modeled as:
@@
 where $\alpha_{t}=1-\beta_{t}$, $\bar{\alpha}_{t}=\prod_{i=1}^{T}\alpha_{i}$, and $T$ represents the total number of timesteps.
-Notably, $\epsilon_{\theta}$ is a denoising network parameterized by $\theta$, which takes $x_{t}$ and $t$ as the inputs to predict the noise needed for denoising.
+Notably, $\epsilon_{\theta}$ is a denoising network parameterized by $\theta$, which takes $x_{t}$ and $t$ as inputs to predict the noise needed for denoising.
 In an image generation procedure using $\mathcal{T}$ timesteps, $\epsilon_{\theta}$ must perform inference $\mathcal{T}$ times, which constitutes the majority of computational costs in diffusion models.
-Recently, numerous studies have shown that implementing $\epsilon_{\theta}$ as a transformer often enhances generation quality.
+Recent studies have shown that implementing $\epsilon_{\theta}$ as a transformer often improves the quality of generation.
 Figure 3: The overview of DuCa.
@@
 Then, DuCa performs a conservative caching step followed by an aggressive caching step, and then another conservative caching step, alternating in this manner to form a complete cache cycle.
-(b) The proposed V-Caching method selects important tokens for DuCa by first fetching the value matrix for all input tokens that have been previously computed in the attention layers and then computing the norm of the value matrix for each token.
-The tokens with the largest V-norm values are selected as cached tokens.
+(b) The proposed method selects tokens for DuCa by setting random scores for all tokens to diversity.
+The tokens with the largest random scores are selected as cached tokens.
 The computation results for these tokens are directly retrieved from the cache, while the remaining tokens are computed through the neural layer, with their updated results stored back into the cache.
@@
 $\mathcal{G}$, $g_{l}$, and $\mathcal{F}^{l}$ represent the DiT model, the blocks in DiT, and different layers in a DiT block, respectively.
-$l$ denotes the index of DiT blocks and $L$ denotes the depth of DiT.
+$l$ denotes the index of DiT blocks, and $L$ denotes the depth of DiT.
 In diffusion transformers, the input data $\mathbf{x}_{t}$ consists of a sequence of tokens representing different patches in the generated images, expressed as $\mathbf{x}_{t}=\{x_{i}\}_{i=1}^{H\times W}$, where $H$ and $W$ indicate the height and width of the images or their latent code, respectively.
 Then, the computation between $\mathcal{F}$ and $\mathbf{x}$ can be generally written as $\mathcal{F}(\mathbf{x})=\mathbf{x}+\text{AdaLN}\circ f(\mathbf{x})$, where the first item $\mathbf{x}$ indicates the residual connection and $f$ denotes the MLP or attention layers.
-AdaLN indicates the adaptive layer normalization [13].
+AdaLN indicates the adaptive layer normalization [Guo2022AdaLNAV].
 Feature Caching: Feature caching aims to reuse the features that are computed and cached in the past timesteps to skip the computation in future timesteps.
 It is usually performed in multiple repeated caching cycles.
-Assuming a caching cycle with $N$ timesteps from the timestep $t$ to $t+N-1$, then feature caching performs the full computation in the timestep $t$ and stores the obtained features in a cache.
+Assuming a caching cycle with $N$ timesteps from the timestep $t$ to t+N1t+N-1, then feature caching performs the full computation in the timestep $t$ and stores the obtained features in a cache.
 This progress can be denoted as $\mathcal{C}[l]:=g_{l}(x_{t})$, where $\mathcal{C}$ denotes the cache and “:=” denotes assigning value operation.
-Then, in timesteps from $t+1$ to $t+N-1$, the computation in these steps is skipped by reusing the cached features, which can be formulated as $g_{l}(x_{t+i}):=\mathcal{C}[l],i\leq N-1$.
-For simplicity, in this paper, the timestep $t$ and the timesteps in $[t+1,t+N-1]$ are named the freshing step and the caching steps, respectively.
-3.2 Dual Feature Caching
+Then, in timesteps from $t+1$ to t+N1t+N-1, the computation in these steps is skipped by reusing the cached features, which can be formulated as $g_{l}(x_{t+i}):=\mathcal{C}[l],i\leq N-1$.
+For simplicity, in this paper, the timestep $t$ and the timesteps in [t+1,t+N1][t+1,t+N-1] are named the freshing step and the caching steps, respectively.
+IV-B Dual Feature Caching
 The proposed dual feature caching aims to perform aggressive caching and conservative caching alternatively in the caching steps.
@@
 Then, the computation with aggressive caching can be written as:
-$\mathcal{G}(x)_{t}:=g_{l+1}\circ...\circ g_{L}^{t}(\mathcal{C}[l]).$ (3)
-To achieve the maximal acceleration with aggressive caching, we set $l=L-1$ in most of our experiments, which indicates almost skipping all the layers in the caching step.
+$\mathcal{G}(x)_{t}:=g_{l+1}\circ...\circ g_{L}^{t}(\mathcal{C}[l]).\vskip-8.53581pt$ (3)
+To achieve the maximal acceleration with aggressive caching, we set l=L1l=L-1 in most of our experiments, which indicates almost skipping all the layers in the caching step.
 Conservative Caching in DuCa
@@
 Instead, it caches and reuses only the features without the residual connection, which allows it to allocate the computation to some layers in the caching steps to correct the cached features.
-Recently, ToCa [35] introduced a token-wise approach that selectively reuses features based on the importance of each token and its sensitivity to caching, which enables us to perform computation only in the important tokens.
+Recently, ToCa [zou2024accelerating] introduced a token-wise approach that selectively reuses features based on the importance of each token and its sensitivity to caching, which enables us to perform computation only in the important tokens.
 Following ToCa, we first calculate an importance score $\mathcal{S}$ for each token $x_{i}$ and then use this score to divide the index set of tokens $\mathcal{I}$ into two complementary sets: $\mathcal{I}_{\text{Cache}}$, for tokens to be cached, and $\mathcal{I}_{\text{Compute}}$, for tokens to be computed.
-Subsequently, the computation for the $i_{th}$ token $x_{i}$ on the $l_{th}$ layer can be expressed as:
-$\gamma_{i}f^{l}(x_{i})+(1-\gamma_{i})\mathcal{C}[l](x_{i}),$ (4)
-where $\gamma_{i}=0$ for $i\in\mathcal{I}_{\text{Cache}}$ and $\gamma_{i}=1$ for $i\in\mathcal{I}_{\text{Compute}}$.
+Subsequently, the computation for the $i_{th}$ token $x_{i}$ on the $l_{th}$ layer can be expressed as: $\gamma_{i}f^{l}(x_{i})+(1-\gamma_{i})\mathcal{C}[l](x_{i}),$ where $\gamma_{i}=0$ for $i\in\mathcal{I}_{\text{Cache}}$ and $\gamma_{i}=1$ for $i\in\mathcal{I}_{\text{Compute}}$.
 Here, $\mathcal{C}[l](x_{i})$ denotes retrieving the cached value of $x_{i}$ from the conservative cache $\mathcal{C}$ at layer $l$, which incurs no computation cost, enabling partial omission of calculations.
 After this selective computation in the token dimension, the computed features are updated to $\mathcal{C}[l]$ for future reusing.
-Value-based Token-wise Feature Caching:
+Key Issue of Attention-Based Selection Methods
 As the previous token-wise conservative caching method, ToCa offers two attention map-based selection strategies that have shown good effectiveness.
-However, the attention map-based selection strategy is incompatible with another commonly used attention-optimization method FlashAttention[9] and memory-efficient attention, which can not provide the attention scores but are capable of reducing the memory costs of attention from $O(N^{2})$ to $O(N)$ and significant accelerating the attention layers.
+However, the attention map-based selection strategy is incompatible with another commonly used attention-optimization method FlashAttention[Dao2022FlashAttentionFA] and memory-efficient attention, which can not provide the attention scores but are capable of reducing the memory costs of attention from $O(N^{2})$ to $O(N)$ and significant accelerating the attention layers.
 This conflict between attention scores-based token selection with this optimized attention computation makes ToCa less practical in application and raises the requirements for attention score-free token selection methods.
-Figure 4 shows the relation between the norm of the value matrix and the attention scores of each token, which demonstrates that tokens with high attention scores in the self-attention map tend to have relatively smaller norms in their value matrix.
-This suggests that we can directly select tokens with smaller norms in their value matrix as substitutes for those with high attention map scores.
-Concretely, the value matrix of each token is produced by a linear layer in attention layers as follows: $v_{i}=\text{MLP}_{v}(x_{i})$, where $v_{i}$ is value vector for the $i^{th}$ token.
-Therefore, V-Caching gives the selection rule for $\mathcal{I}_{\text{Cache}}$ as follows:
-$\mathcal{I}_{\text{Cache}}=\mathop{\arg\max}_{\{i_{1},i_{2},\dots,i_{n}\}\subseteq\{1,2,\dots,N\}}\{\|v_{i_{1}}\|_{2},\|v_{i_{2}}\|_{2},\cdots,\|v_{i_{n}}\|_{2}\},$ (5)
-where the $n$ is the total number of selected tokens, defined by $n=\lfloor R\%\times N\rfloor$.
-3.3 Overall Framework
-Built on the deeper understanding of aggressive and conservative caching, DuCa introduces a caching scheme that alternates conservative and aggressive caching in different steps.
-V-Caching serves as the conservative caching strategy, as shown in Figure 3.
+TABLE I: Token selection methods for FLUX on Image Reward.
+The experiments were conducted on the FLUX.1-dev model with a full activation interval of 4, selecting $10\%$ of tokens at each cache step.
+A higher Image Reward means better performance.
+Score Efficient Attention Max Min
+Attention 0.9798 0.9707
+K-norm 0.9783 0.9726
+V-norm 0.9747 0.9798
+Similarity 0.9715 0.9811
+Random 0.9806
+Original Model 0.9898
+Diversity matters more in token selection
+As shown in Table I, we compare attention scores with other classic token selection methods used in large language and multimodal models, such as K-norm and V-norm, which select tokens based on maximizing and minimizing the corresponding values, respectively.
+We then compare these results with the simplest and most naive random selection.
+Surprisingly, while the method of selecting tokens based on the highest attention score in ToCa is the best-performing approach among all standard-based selection methods, it still performs slightly worse than random selection.
+The results suggest that rather than the traditional assumption that tokens can be distinguished as important or unimportant, selecting a diverse and broad range of tokens is more beneficial for performance.
+Based on this hypothesis, we designed an experiment that scores tokens based on similarity.
+For each cache step in Table I, $10\%$ of all tokens were selected.
+First, $1\%$ of all tokens were randomly chosen as the base tokens.
+Then, for the remaining $9\%$, tokens were selected based on two criteria: one method selects tokens with the highest similarity to the base tokens, aiming to choose tokens of the same type, while the other method selects those with the lowest similarity, encouraging greater diversity in the selected tokens.
+The method based on maximum similarity achieved the worst performance among all selection methods, while the method based on minimum similarity achieved the best performance, further confirming the importance of diversity in token selection.
+Considering that the difference between the minimum similarity-based method and the random selection method is small and it does not require additional computation, the random token selection method is recommended.
+IV-C Overall Framework
+Based on the understanding of aggressive and conservative caching, DuCa introduces a caching scheme that alternates conservative and aggressive caching in different steps.
+Random selection serves as the selection strategy for conservative caching, as shown in Figure 3.
 This design leverages the realignment capability of token-wise conservative caching to correct the misaligned timestep information occurring in aggressive caching steps.
 Therefore, it enables the high local compression rate of aggressive caching to be applied to more steps while further reducing redundant computations in conservative caching steps.
-In practice, we arrange token-wise conservative caching steps on odd-numbered steps following a fresh step, and aggressive caching steps on even-numbered steps.
-Figure 4: The relation between the self-attention score and the norm of the value matrix for each token. Tokens with the largest attention score tend to have a lower norm in the value matrix.
-4 Experiments
-4.1 Experiment Settings
-4.1.1 Model Configurations
-Experiments are conducted on three widely known DiT-based models on the three generation tasks with the correlating default samplers, including PixArt-$\alpha$ [7] with 20 DPM Solver++ [23] steps for text-to-image generation, OpenSora [34] with 30 reflected flow steps [21] for text-to-video generation, and DiT-XL/2 [26] with 50 DDIM [31] steps for class-conditional image generation.
-All the experiments are conducted with NVIDIA H800 80GB GPUs.
-See More details in our appendix.
-4.1.2 Evaluation and Metrics
-Text-to-image generation We randomly select 30,000 captions from COCO-2017 datasets [20] to generate the 30,000 $256\times 256$ images.
-FID-30k is computed using the 30k generated images and the real images from COCO-2017 to evaluate the quality of text-to-image generation.
-Besides, the CLIP score [14] is computed between the 30k prompt-image pairs to evaluate the text-image alignment.
-All the CLIP scores are tested using the Vit-large-14 model.
-Text-to-video generation VBench [17] evaluation framework is applied to evaluate the performance of the text-to-video model.
-With 5 videos per prompt and 950 prompts from the framework, 4,750 videos are generated to evaluate the 16 aspects proposed by VBench.
-The generated videos are 480p, 9:16, and 2s videos.
-Class-conditional image generation 50,000 $256\times 256$ images uniformly sampled from the 1,000 classes in ImageNet[10] are generated to evaluate the FID-50k [15].
-Besides, sFID, Precision, and Recall are introduced as supplementary metrics.
-Table 1: Quantitative comparison of text-to-image generation on MS-COC02017 with PixArt-$\alpha$ and 20 DPM++ sampling steps.
+We arrange token-wise conservative caching steps on odd-numbered steps following a fresh step and aggressive caching steps on even-numbered steps.
+TABLE II: Quantitative comparison in text-to-image generation for FLUX on Image Reward.
+[…]
+FLUX.1[flux2024] Attention [Dao2022FlashAttentionFA] Latency(s) $\downarrow$ Speed $\uparrow$ FLOPs(T) $\downarrow$ Speed $\uparrow$ DrawBench Overall
 […]
 […]
-PixArt-$\alpha$ [7] 0.242 11.88 1.00$\times$ 28.12 16.29
-$50\%$ steps 0.121 5.94 2.00$\times$ 37.57 15.83
-FORA [28] 0.178 5.66 1.98$\times$ 29.67 16.40
-ToCa[35] 0.167 6.05 1.96$\times$ 28.38 16.43
-DuCa(C-A score) 0.157 5.99 1.98$\times$ 27.98 16.44
-DuCa(V-Caching) 0.088 5.99 1.98$\times$ 28.05 16.42
-Table 2: Quantitative comparison in text-to-video generation on VBench. *Results are from PAB [33]. PAB1-3 indicate PAB with different hyper-parameters.
-[…]
-OpenSora [34] 54.83 3283.20 1.00$\times$ 79.41
 […]
 […]
-PAB1∗ [33] 41.05 2657.70 1.24$\times$ 78.51
-PAB2∗ [33] 39.96 2615.15 1.26$\times$ 77.64
-PAB3∗ [33] 38.26 2558.25 1.28$\times$ 76.95
-$50\%$ steps 28.85 1641.60 2.00$\times$ 76.78
-FORA [28] 33.19 1751.32 1.87$\times$ 76.91
-ToCa[35] 29.03 1394.03 2.36$\times$ 78.59
-DuCa(C-A score) 26.77 1315.62 2.50$\times$ 78.63
-DuCa(V-Caching) 19.16 1315.62 2.50$\times$ 78.54
-Table 3: Quantitative comparison on class-to-image generation on ImageNet with DiT-XL/2.
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+V Experiments
+V-A Experiment Settings
+V-A1 Model Configurations
+Experiments are conducted on four widely known DiT-based models on the three generation tasks with the correlating default samplers, including the FLUX.1-dev model [flux2024] with 50 rectified flow sampling steps, HiDream model with 50 rectified flow sampling steps, OpenSora [opensora] with 30 rectified flow sampling steps [refitiedflow] for text-to-video generation, and DiT-XL/2 [peebles2023dit] with 50 DDIM [songDDIM] steps for class-conditional image generation.
+All the experiments are conducted with NVIDIA H800 80GB GPUs.
+Additionally, we have supplemented our experiments with the smaller text-to-image generation model PixArt-$\alpha$ [chen2023pixartalpha] with 20 DPM Solver++ [lu2022dpm++] steps and included the detailed design of all experiments.
+More details are available in the appendix.
+V-A2 Evaluation and Metrics
+Text-to-image generation For the FLUX.1-dev and HiDream model with a high resolution of $1000\times 1000$, we utilized 200 high-quality prompts from DrawBench [Saharia2022PhotorealisticTD], which focus on generating various types of images.
+We then generated images using these prompts and employed Image Reward [xuImageRewardLearningEvaluating2023] to assess the quality of the generated images while using the GenEval [Ghosh2023GenEvalAO] benchmarks to evaluate the alignment between the images and the text.
+Text-to-video generation VBench [huang2024vbench] evaluation framework is applied to evaluate the performance of the text-to-video model.
+With 5 videos per prompt and 950 prompts from the framework, 4,750 480p, 9:16, and 2s videos are generated to evaluate the 16 aspects proposed by VBench.
+TABLE III: Quantitative results for HiDream-l1.
+Method Acceleration Image Reward $\uparrow$
+HiDream-l1 Latency(s) $\downarrow$ Speed $\uparrow$ FLOPs(T) $\downarrow$ Speed $\uparrow$ DrawBench
+[full]: 50 steps 48.81 $\times 1.00$ 7780.0 $\times 1.00$ 1.182
+25 steps 24.41 $\times 2.00$ 3890.0 $\times 2.00$ 1.128
+ToCa ($N=4$) 31.22 $\times 1.56$ 2540.1 $\times 3.06$ 1.111
+PAB ($N=8$) 36.33 $\times 1.34$ 5962.0 $\times 1.30$ 0.8876
+DuCa ($N=3,R=90\%$) 19.88 $\times 2.46$ 2873.4 $\times 2.71$ 1.152
+DuCa ($N=4,R=90\%$) 15.36 $\times 3.18$ 2197.3 $\times 3.54$ 1.090
+Class-conditional image generation 50,000 $256\times 256$ images uniformly sampled from the 1,000 classes in ImageNet[deng2009imagenet] are generated to evaluate the FID-50k [heusel2017fid].
+Besides, sFID, Precision, and Recall are introduced as supplementary metrics.
+V-A3 Results on text-to-image generation
+The generation results of the proposed DuCa under various configurations are compared against recent state-of-the-art acceleration methods, including ToCa [zou2024accelerating], FORA [selvaraju2024fora], and a series of methods that directly reduce sampling steps.
+These comparisons are conducted on FLUX.1-dev [flux2024] and HiDream at a resolution of $1000\times 1000$, as detailed in Table II and III.
+[…]
+[…]
+TABLE IV: Quantitative comparison on class-to-image generation on ImageNet with DiT-XL/2.
 Method Latency(s) $\downarrow$ FLOPs(T) $\downarrow$ Speed $\uparrow$ sFID $\downarrow$ FID $\downarrow$ Precision $\uparrow$ Recall $\uparrow$
-DiT-XL/2-G (cfg $=1.50$) 2.012 118.68 1.00$\times$ 4.98 2.31 0.82 0.58
-DDIM-50 steps 0.417 23.74 1.00$\times$ 4.40 2.43 0.80 0.59
-DDIM-25 steps 0.209 11.87 2.00$\times$ 4.74 3.18 0.79 0.58
-DDIM-20 steps 0.167 9.49 2.50$\times$ 5.15 3.81 0.78 0.58
-DDIM-17 steps 0.142 8.07 2.94$\times$ 5.76 4.58 0.77 0.56
-FORA [28] 0.197 4.29 2.77$\times$ 6.43 3.88 0.79 0.54
-ToCa(a) 0.195 8.73 2.72$\times$ 5.14 3.64 0.79 0.55
-ToCa(b) 0.214 10.23 2.32$\times$ 4.70 3.05 0.79 0.57
-DuCa(a)(S-A score) 0.184 8.76 2.71$\times$ 4.66 3.19 0.79 0.57
-DuCa(b)(S-A score) 0.200 9.58 2.48$\times$ 4.66 3.05 0.80 0.57
-DuCa(a)(V-Caching) 0.177 8.76 2.71$\times$ 4.68 3.19 0.80 0.57
-DuCa(b)(V-Caching) 0.198 9.58 2.48$\times$ 4.69 3.06 0.80 0.57
-4.2 Quantitative and Qualitative Results
-Figure 5: Visualization results for different acceleration methods on PixArt-$\alpha$.
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+DuCa(b) 0.198 $8×\times 4.59 2$.84 0.80 0.58
+V-B Quantitative and Qualitative Results
+Figure 4: Visualization results for different acceleration methods on FLUX.1-dev model.
+[…]
+[…]
+V-B1 Results on text-to-video generation
+The generation results of the state-of-the-art acceleration methods $\Delta$-DiT, T-GATE[tgatev1], PAB[zhao2024PAB], 15 rflow sampling steps, ToCa[zou2024accelerating], FORA [selvaraju2024fora] and proposed DuCa on the text-to-video model OpenSora[opensora] are shown in Table V.
+As shown in Table V, DuCa achieves a further acceleration from 2.36 $\times$ (as with ToCa) to 2.50 $\times$ with virtually no loss in generation quality.
+Additionally, despite having the lowest computational cost among all the compared methods, DuCa still achieves the highest VBench score of 78.83, demonstrating that the proposed strategy in DuCa, which alternates between aggressive and conservative caching steps, effectively eliminates the redundant computations in the fully conservative caching approach.
 Figure 5: Visualization results for different methods on OpenSora (videos available in the supplementary material).
-4.2.1 Results on text-to-image generation
-As shown in Table 1, we compared the DuCa under two different token selection methods: Cross-Attention score from ToCa and the proposed V-norm score with three training-free acceleration methods: ToCa[35], the 10 DPM++ sampling steps and FORA [28], a layer-wise conservative caching method.
-In terms of generation quality, the quantitative results demonstrate that under an acceleration ratio close to 2.0, the DuCa methods achieve the lowest FID-30k score on the MS-COCO2017 generation task, with almost no loss in generation quality compared to the original non-accelerated approach.
-For instance, the DuCa method using the Cross-Attention score reduces the FID-30k by 0.40 compared to the ToCa method with the same acceleration ratio, while the DuCa method using the V-norm score also decreases the FID-30k by 0.33, demonstrating that our alternating acceleration strategy effectively enhances generation quality.
-In terms of evaluating the text-image alignment metric, the performance of the DuCa method is very close to that of ToCa and generally superior to the approach of directly setting $50\%$ generation steps.
-For example, the DuCa method with Cross-Attention score and V-norm score achieves CLIP scores on MS-COCO2017 that are 0.61 and 0.59 higher, respectively, compared to the $50\%$ generation steps method with the same acceleration ratio.
-These scores are even slightly better than those of the original non-accelerated PixArt-$\alpha$ model, with improvements of 0.15 and 0.13, respectively, demonstrating better controllability.
-In addition, thanks to the acceleration brought by the compatibility with FlashAttention [9], the latency of the DuCa method with V-norm score is reduced by 0.069 compared to the DuCa method with Attention score, representing an approximate $43.9\%$ reduction, which is a significant improvement in acceleration.
-4.2.2 Results on text-to-video generation
-Figure 7: VBench metrics and acceleration ratio of proposed DuCa and other methods
-The generation results of $\Delta$-DiT, T-GATE[32], PAB[33], 15 rflow sampling steps, ToCa[35], FORA [28] and proposed DuCa on the text-to-video model OpenSora[34] are shown in Table 2.
-As shown in Table 2, DuCa achieves a further acceleration from 2.36 $\times$ (as with ToCa) to 2.50 $\times$ with virtually no loss in generation quality.
-Additionally, despite having the lowest computational cost among all the compared methods, DuCa still achieves the highest VBench score.
-This demonstrates that the proposed strategy in DuCa, which alternates between aggressive and conservative caching steps, effectively eliminates the redundant components present in a fully conservative caching approach.
-Additionally, comparing the generation results of the Cross-Attention and V-Caching token selection strategies, we only observe a slight decrease of 0.09 in the VBench score with the V-Caching strategy.
-More significantly, V-Caching reduces generation time by 7.61 seconds, approximately $28.4\%$.
-This indicates that DuCa effectively leverages V-Caching to make the acceleration method compatible with FlashAttention[9], achieving substantial reductions in generation time.
-In practice, using V-Caching will be more advantageous for rapid inference.
+TABLE V: Quantitative comparison in text-to-video generation on VBench. *Results are from PAB [zhao2024PAB]. PAB1-3 indicate PAB with different hyper-parameters.
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+The advantages of DuCa, stemming from its diverse selection strategies, are also evident in its comparison with ToCa: although DuCa eliminates unnecessary redundant computations, resulting in lower actual computational load than ToCa, its VBench score decreases by only 0.58 compared to the original model, while ToCa’s score drops by 0.82, indicating that DuCa reduces quality loss by $29.3\%$.
+Furthermore, thanks to DuCa’s compatibility with FlashAttention [Dao2022FlashAttentionFA], the generation of a single 480p, 2s video takes only 19.16s with DuCa, compared to 29.03s with ToCa, resulting in a speedup of nearly 10s.
 In Figure 6, we further present the performance of DuCa and the aforementioned methods across various dimensions of the VBench metric.
-DuCa shows results consistent with the non-accelerated original OpenSora model across multiple metrics, including but not limited to Overall Consistency, Temporal Style, Scene, and Object Class.
-Additionally, the Cross-Attention and V-Caching strategies introduce only a slight difference in Spatial Relationship, further underscoring the exceptional performance of DuCa and the proposed V-Caching in practical scenarios.
-4.2.3 Results on class-conditional image generation
-In Table 3, we compare ToCa[35], FORA [28], the direct reduction of corresponding DDIM steps, and the proposed DuCa under Self-Attention score and V-Caching token selection strategies.
+DuCa shows results consistent with the non-accelerated original OpenSora model across multiple metrics.
+As shown in Figure 5, it is evident that DuCa is closer to the original model compared to other methods.
+V-B2 Results on class-conditional image generation
+In Table IV, we compare ToCa[zou2024accelerating], FORA [selvaraju2024fora], the direct reduction of the corresponding DDIM steps and the proposed DuCa.
 The experimental results indicate that with comparable FID scores, representing generation quality, DuCa further enhances acceleration.
-For instance, DuCa(b) with Self-Attention score achieves a similar FID-50k as ToCa(b), but the acceleration ratio increases from 2.36 to 2.48.
-Additionally, under a higher acceleration scenario with an acceleration ratio around 2.7, DuCa achieves an FID-50k of 3.19, further reducing the FID-50k by 0.45 compared to ToCa’s 3.64.
-This demonstrates that DuCa effectively reallocates computation to more critical areas.
-4.3 Ablation Studies
-Table 4: Ablation Study for different token selection methods for DuCa on PixArt-$\alpha$.
-Methods MS-COCO2017 Latency(s) $\downarrow$ FLOPs(T) $\downarrow$
-FID-30k $\downarrow$ CLIP $\uparrow$
-Random 28.34 16.42 0.089 5.99
-Cross-Attention 27.98 16.44 0.157 5.99
-V-Caching 28.05 16.42 0.088 5.99
-Table 5: Comparison on different caching methods on PixArt-$\alpha$.
-Methods MS-COCO2017 FLOPs(T) $\downarrow$
-FID-30k $\downarrow$ CLIP $\uparrow$
-Conservative-only 27.92 16.39 7.84
-Aggressive-only 30.51 16.39 3.59
-DuCa 28.05 16.42 5.99
-As shown in Table 4, we compare Random, Cross-Attention based, and V-Caching token selection methods on PixArt-$\alpha$ in terms of latency, FLOPs, FID-30k, and CLIP score.
-It can be observed that the cross-Attention-based selection method achieves the best performance in generation quality and text-image alignment, but due to its incompatibility with FlashAttention[9], its latency is as high as 0.157 seconds—nearly twice that of V-Caching and Random selection.
-In contrast, V-Caching achieves nearly the same, only slightly lower, FID-30k and CLIP scores, along with faster generation speed.
-Despite the minimal impact of the three token selection methods on computational load, the FlashAttention[9]-compatible V-Caching method shows significantly higher acceleration due to computational optimizations.
-Thus, V-Caching stands out as a superior generation scheme that effectively balances quality and efficiency.
-As shown in Table 5, we compare the Conservative-only, Aggressive-only, and alternating DuCa structures on PixArt-$\alpha$ in terms of FLOPs, FID-30k, and CLIP score.
-The Conservative-only structure appears to have the best generation quality but incurs an excessively high computational cost.
+For instance, DuCa(b) with Self-Attention score achieves a similar FID-50k as ToCa(b), but the acceleration ratio increases from 2.32 to 2.48.
+Additionally, under a higher acceleration scenario with an acceleration ratio around 2.7, DuCa achieves an FID-50k of 3.00, further reducing the FID-50k by 0.31 compared to ToCa’s 3.31, demonstrating that DuCa effectively reallocates computation.
+Figure 6: VBench metrics and acceleration ratio of proposed DuCa and other methods.
+Figure 7: Ablation study on cycle length NN and cache ratio $R$. Setting $N=5$ and $R=0.9$ yields the best trade-off between computation and quality.
+V-C Ablation Studies
+TABLE VI: Different caching methods on FLUX.1-dev.
+Methods Image Reward FLOPs(T) $\downarrow$
+Conservative-only 0.9789 1263.22
+Aggressive-only 0.8539 1042.12
+DuCa 0.9895 1190.24
+Diversity in Token Selection
+As demonstrated in Table I, the results prove that the adopted random selection method, due to its consideration of diversity in token selection, outperforms a series of traditional methods based on specific metrics while maintaining compatibility with the Efficient Attention framework.
+Additionally, through experiments designed around similarity, we have explored and strongly validated the importance of diversity in token selection.
+As shown in Table VI, we compare the conservative-only, aggressive-only, and alternating DuCa structures, where the aggressive structure is used in activated odd steps and the conservative structure in even steps, on FLUX.1-dev in terms of Image Reward and FLOPs.
+The conservative-only structure appears to have good generation quality but incurs an excessively high computational cost.
 In contrast, aggressive-only reduces computation to nearly half of conservative-only but suffers significantly in generation quality.
-Our proposed DuCa structure achieves a computational load between the two, with only a slight 0.13 increase in FID-30k compared to the high-computation Conservative-only approach, while improving text-image alignment.
-This indicates that DuCa effectively eliminates most redundant computation present in Conservative-only.
-5 Conclusion
-This paper first gives an in-depth study of the caching error on aggressive and conservative methods, and then introduces dual caching DuCa, which performs aggressive and conservative alternatively to leverage their advantages in different caching steps.
-Besides, V-Caching is introduced to perform token-wise conservative caching based on the norm of the value matrix of the tokens, making it compatible with default optimized attention computation such as FlashAttention, allowing feature caching to be practical in real applications.
-We hope this paper can attract more attention to feature caching for its real benefits, to make it simple but effective, instead of being complex.
-Supplementary Material
-6 Experimental Details
-In this section, more details of the experiments are provided.
-6.1 Model Configuration
-As mentioned in 4.1.1 , experiments on three models from different tasks, PixArt-$\alpha$[7] for text-to-image generation, OpenSora[4] for text-to-video generation, and DiT[26] for class-conditional image generation, are presented.
-In this section, a more detailed hyperparameter configuration scheme is provided.
-  • PixArt-$\alpha$: The average cache cycle length is set to $\mathcal{N}=3$ with an average caching ratio of $R=70\%$ for ToCa[35], and $\mathcal{N}=3$ with $\mathcal{R}=25\%$ for the DuCa method to maintain a similar acceleration ratio.
-Additionally, in the FORA [28] scheme, $\mathcal{N}=2$.
-Conservative caching steps are configured to occur on odd-numbered steps following fresh steps, while aggressive caching steps occur on even-numbered steps following fresh steps.
-The proportion of conservative caching steps across layers and steps is identical to that in ToCa[35].
-  • OpenSora: The average cache cycle length is $\mathcal{N}=3$ for temporal attention, spatial attention, and MLP, and $\mathcal{N}=6$ for cross-attention, with an average caching ratio $\mathcal{R}=85\%$ exclusively for the MLP in ToCa[35] and DuCa, while all other modules use a caching ratio of $100\%$.
-In the FORA [28]scheme, all modules are set to $\mathcal{N}=2$.
-    For $\textbf{PAB}^{1*}$ [33], temporal attention is set to $\mathcal{N}=2$, spatial attention to $\mathcal{N}=4$, cross-attention to $\mathcal{N}=6$, and the MLP module is fully computed.
-For $\textbf{PAB}^{2*}$ [33], $\mathcal{N}$ is set to 3, 5, and 7 for temporal, spatial, and cross-attention, respectively.
-For $\textbf{PAB}^{3*}$ [33], $\mathcal{N}$ is set to 5, 7, and 9 for temporal, spatial, and cross-attention, respectively.
-    For DuCa, conservative caching steps are set on odd-numbered steps following fresh steps, while aggressive caching steps are set on even-numbered steps following fresh steps.
-The proportion of conservative caching steps across layers and steps is identical to that in ToCa[35].
-  • DiT: The average cache cycle length is set to $\mathcal{N}=4$ with an average caching ratio of $R=93\%$ for ToCa[35](a), $\mathcal{N}=3$ with $R=93\%$ for ToCa[35](b), and $\mathcal{N}=3$ with $R=95\%$ for both DuCa(a) and (b).
-Additionally, the average cache cycle length for FORA [28]is set to $\mathcal{N}=3$.
-    Furthermore, for DuCa(a), conservative caching steps are set on even-numbered steps following fresh steps, while aggressive caching steps are set on odd-numbered steps.
-In contrast, for DuCa(b), the arrangement is reversed.
-    Since the average cache cycle length $\mathcal{N}=3$ here, DuCa(a), which places aggressive caching steps immediately after the fresh step, achieves a higher acceleration ratio compared to DuCa(b), which places conservative caching steps in the first position after the fresh step.
-6.2 Anonymous Page for video presentation
-To further demonstrate the superiority of DuCa in video generation, we have created an anonymous GitHub page.
-Please visit https://duca2024.github.io/DuCa/ for a more detailed view.
-Besides, the videos are also available through the Supplementary Material.
+Our proposed DuCa structure achieves a balanced computational load between the two, reducing the computational cost compared to the conservative-only method while maintaining comparable or even better generation quality.
+Choice of Cycle Length and Cache Ratio
+We further conduct ablation experiments to investigate the influence of cycle length ($N$) and cache ratio ($R$) on performance, as illustrated in Figure 7.
+When varying $N$ from 3 to 8, we observe that shorter cycles incur higher FLOPs, while excessively long cycles result in noticeable quality degradation.
+The setting $N=5$ provides a favorable trade-off, preserving generation quality while reducing computation.
+Similarly, for cache ratio $R$, lowering $R$ below $0.7$ significantly harms quality, whereas $R=0.9$ eliminates redundant computation without compromising fidelity.
+Therefore, we adopt $N=5$ and $R=0.9$ as default hyperparameters in our main experiments.
+VI Conclusion
+This paper first gives an in-depth study of the caching error on aggressive and conservative methods and then introduces dual caching DuCa, which performs aggressive and conservative alternatively to leverage their advantages in different caching steps.
+Moreover, we have discovered the value of random selection and further explored its implications, revealing that diversity plays a more critical role in token selection than previously thought.
+Random selection not only demonstrates excellent performance but also seamlessly integrates with efficient attention methods like FlashAttention, making feature caching a practical and viable approach in real-world applications.
+We hope this paper can attract more attention to feature caching for its real benefits, to make it simple but effective, instead of complex.